WHAT MASS OF ZINC CAN BE PRODUCED BY THE ELECTROLYSIS OF ZINC SULPHATE SOLUTION WHEN STEADY CURRENT OF 0.015A IS PASSED FOR 15 MINUTES.


Grade:11

2 Answers

upender surepally
10 years ago

0.049GRAMS

Manas Satish Bedmutha
10 years ago
mass(grams) = Z * charge(amp-hr) =1.22*0.015*15/60 =1.22*0.015*0.25 =0.4575g